OpenStudy (anonymous):

PLEASE HELP!! I GIVE MEDAL If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 15x - y = -6 5x - 3y = -13

OpenStudy (anonymous):

so first you want to rewrite the first equation so that you get it into y = x + b. So you want to subtract 15x from each side so that you get -y= -15x - 6

OpenStudy (anonymous):

No do the same thing for the second equation and you get -3y = -5x - 13

OpenStudy (anonymous):

you want make the first equation y instead of -y so divide by negative one. Now it is y = 15x + 6. So substitute in 15x + 6 into the second equation so that you get -3(15x+6) = -5x - 13

OpenStudy (anonymous):

now all that's left to do is to solve the final equation

OpenStudy (anonymous):

So distribute \[-3(15x) + -3(6); -45x - 18 \]

OpenStudy (anonymous):

\[-45x-18 = -5x - 13\]

OpenStudy (anonymous):

add 5x to each side and add 18 to each side and you get\[-50x = 5\]

OpenStudy (anonymous):

all that's left to do is simplify by dividing -50 on each side and you get x = -0.1

OpenStudy (anonymous):

:) @jackmullen55